# https://uni-fl.github.io/

### v1

@@ Abstract
Diffusion models have revolutionized the field of image generation, leading to the proliferation of high-quality models and diverse downstream applications. However, despite these significant advancements, the current competitive solutions still suffer from several limitations, including inferior visual quality, a lack of aesthetic appeal, and inefficient inference, without a comprehensive solution in sight. To address these challenges, we present UniFL, a unified framework that leverages feedback learning to enhance diffusion models comprehensively. UniFL stands out as a universal, effective, and generalizable solution applicable to various diffusion models, such as SD1.5 and SDXL. Notably, UniFL incorporates three key components: perceptual feedback learning, which enhances visual quality; decoupled feedback learning, which improves aesthetic appeal; and adversarial feedback learning, which optimizes inference speed. In-depth experiments and extensive user studies validate the superior performance of our proposed method in enhancing both the quality of generated models and their acceleration. For instance, UniFL surpasses ImageReward by 17% user preference in terms of generation quality and outperforms LCM and SDXL Turbo by 57% and 20% in 4-step inference. Moreover, we have verified the efficacy of our approach in downstream tasks, including Lora, ControlNet, and AnimateDiff.
Keywords:
Diffusion Models Feedback Learning Acceleration
Figure 1: We propose UniFL, a unified framework that leverages feedback learning to elevate the visual generation quality, enhance preference aesthetics, and accelerate the inference process. The figure illustrates the outcomes obtained by optimizing SDXL through UniFL, and the last three images depict the results with 4 steps inference.
@@ 1 Introduction
The emergence of diffusion models has catapulted the field of Text-to-Image (T2I) into a realm of unparalleled progress, marked by notable contributions like DALLE-3 [34], Imagen [38], Midjourney [49], and etc,. In particular, the introduction of open-source image generation models, exemplified by stable diffusion [35], has inaugurated a transformative era of Text-to-Image, giving rise to numerous downstream applications such as T2I Personalization [37, 15, 21, 60], Controllable Generation [61, 33, 29] and Text-to-Video (T2V) Generation [53, 16, 18]. Despite the remarkable progress achieved thus far, current stable diffusion-based image generation models still exhibit certain limitations. i) Inferior Quality: the generated images often suffer from poor quality and lack authenticity. Examples include characters with incomplete limbs or distorted body parts, as well as limited fidelity in terms of style representation. ii) Lack Aesthetics: there is a notable bias in the aesthetic appeal of the generated images, as they often fail to align with human preferences. Deficiencies in crucial aspects such as details, lighting, and atmosphere further contribute to this aesthetic disparity. iii) Inference Inefficiency: the iterative denoising process employed by diffusion models introduces inefficiencies that significantly impede inference speed, thereby limiting the practicality of these models in various application scenarios.
@@
Recently, numerous works have endeavored to address the aforementioned challenges. For instance, SDXL [32] enhances the generation quality of diffusion models by refining training strategies, while RAPHAEL [59] resorts to the techniques of Mixture of Experts (MoE) [44, 63, 14]. RAFT [11], HPS [55, 54], ImageReward [57], and DPO [50] propose incorporating human feedback to guide diffusion models toward aligning with human preferences. SDXL Turbo [40], PGD [39], and LCM [27, 28], on the other hand, tackle the issue of inference acceleration through techniques like distillation and consistency models [46]. However, these methods primarily concentrate on tackling individual problems through specialized designs, which poses a significant challenge for the direct integration of these techniques. For example, MoE significantly complicates the pipeline, making the acceleration method infeasible, and the consistency models [46] alter the denoising process of the diffusion model, making it arduous to directly apply the ReFL framework proposed by ImageReward [57]. Naturally, a pertinent question arises: Can we devise a more effective approach that comprehensively enhances diffusion models in terms of image quality, aesthetic appearance, and generation speed?
In this paper, we present UniFL, an innovative approach that offers a comprehensive improvement to diffusion models through unified feedback learning. UniFL aims to elevate the visual generation quality, enhance preference aesthetics, and accelerate the inference process. To achieve these objectives, we propose three key components. Firstly, we introduce a pioneering perceptual feedback learning (PeFL) framework that effectively harnesses the extensive knowledge embedded within diverse existing perceptual models to improve visual generation quality. This framework enables the provision of more precise and targeted feedback signals, ultimately enhancing the quality of visual generation in various aspects.
Secondly, we employ decoupled feedback learning to optimize aesthetic quality. By breaking down the coarse aesthetic concept into distinct aspects such as color, atmosphere, and texture, UniFL simplifies the challenge of aesthetic optimization. Additionally, we introduce an active prompt selection strategy to choose the prompts that are more informative and diverse to facilitate more efficient aesthetics preference feedback learning.
Lastly, UniFL develops adversarial feedback learning, wherein the reward model and diffusion model are trained adversarially, enabling the samples under the low denoising steps to be well optimized via the reward feedback, and finally achieves superior inference acceleration. UniFL presents a unified formulation of feedback learning that is both straightforward and versatile, making it adaptable to a wide range of models and yielding impressive improvements. Our contributions are summarized as follows:
@@
  • Novelty and Pioneering: In our work, we shed light on the untapped potential of leveraging existing perceptual models in feedback learning for diffusion models. We highlight the significance of decoupled reward models and elucidate the underlying acceleration mechanism through adversarial training. We believe our ablation experiments provide valuable insights that enrich the community’s understanding of these techniques.
  • High Effectiveness: Through extensive experiments, we demonstrate the substantial improvements achieved by UniFL across multiple types of diffusion models, including SD1.5 and SDXL, in terms of generation quality and acceleration. Furthermore, UniFL outperforms competitive existing approaches and exhibits strong generalization on various downstream tasks.
2 Related Work
2.1 Text-to-Image Diffusion Models
Recently, diffusion models have gained substantial attention and emerged as the de facto method for text-to-image generation, surpassing traditional probabilistic models like GAN [17] and VAE [22]. Numerous related works have been proposed, including GLIDE [30], DALL-E2 [34], Imagen [38], CogView [10] and many others. Among these, Latent Diffusion Models (LDM) [35] extend the diffusion process to the latent space and significantly improve the training and inference efficiency of the diffusion models, opening the door to diverse applications such as controllable generation [61, 33], image editing [29, 19, 3], and image personalization [37, 21, 15] and so on. Despite the progress achieved thus far, current text-to-image diffusion models still have limitations in inferior visual generation quality, deviations from human aesthetic preferences, and inefficient inference. The target of this work is to offer a comprehensive solution that effectively addresses these issues.
2.2 Improvements on Text-to-Image Diffusion Models
Given the aforementioned limitations, researchers have proposed various methods to tackle these issues. Notably,  [32, 59, 6] focuses on improving generation quality through more advanced training strategies. Building upon the success of reinforcement learning with human feedback (RLHF) [31, 1] in the field of LLM,  [55, 54, 2, 57, 64] explore the incorporation of human feedback to improve image aesthetic quality. On the other hand,  [39, 41, 46, 27, 28] concentrate on acceleration techniques, such as distillation and consistency models [46] to achieve inference acceleration. While these methods have demonstrated their effectiveness in addressing specific challenges, their independent nature makes it challenging to combine them for comprehensive improvements. In contrast, our study unifies the objective of enhancing visual quality, aligning with human aesthetic preferences, and acceleration through the feedback learning framework.
@@ 3 Preliminaries
Text-to-Image Diffusion Model. Text-to-image diffusion models leverage diffusion modeling to generate high-quality images based on textual prompts via the diffusion model, which generates desired data samples from Gaussian noise through a gradual denoising process. During pre-training, a sampled image $x$ is first processed by a pre-trained VAE encoder to derive its latent representation $z$. Subsequently, random noise is injected into the latent representation through a forward diffusion process, following a predefined schedule $\{\beta_{t}\}^{T}$. This process can be formulated as $z_{t}=\sqrt{\overline{{\alpha}}_{t}}z+\sqrt{1-\overline{{\alpha}}_{t}}\epsilon$, where $\epsilon\in\mathcal{N}(0,1)$ is the random noise with identical dimension to $z$, $\overline{{\alpha}}_{t}=\prod_{s=1}^{t}\alpha_{s}$ and $\alpha_{t}=1-\beta_{t}$. To achieve the denoising process, a UNet $\epsilon_{\theta}$ is trained to predict the added noise in the forward diffusion process, conditioned on the noised latent and the text prompt $c$. Formally, the optimization objective of the UNet is:
@@
Reward Feedback Learning. Reward feedback learning(ReFL) [57] is a preference fine-tuning framework that aims to improve the diffusion model via human preference feedback. It primarily includes two phases: (1) Reward Model Training and (2) Preference Fine-tuning. In the Reward Model Training phase, human preference data is collected. These data are then utilized to train a human preference reward model, which serves as an encoding mechanism for capturing human preferences. More specifically, considering two candidate generations, denoted as $x_{w}$ (preferred generation) and $x_{l}$ (unpreferred one), the loss function for training the human preference reward model $r_{\theta}$ can be formulated as follows:
$\mathcal{L}(\theta)_{rm}=-\mathbb{E}_{(c,x_{w},x_{l})\sim\mathcal{D}}[log(\sigma(r_{\theta}(c,x_{w})-r_{\theta}(c,x_{l})))]$ (2)
where $\mathcal{D}$ denotes the collected feedback data, $\sigma(\cdot)$ represents the sigmoid function, and $c$ corresponds to the text prompt. The reward model $r_{\theta}$ is optimized to produce a preference-aligned score that aligns with human preferences. In the Preference Fine-tuning phase, ReFL begins with an input prompt $c$, initializing a latent variable $x_{T}$ at random. The latent variable is then progressively denoised until reaching a randomly selected timestep $t$. At this point, the denoised image $x^{\prime}_{0}$ is directly predicted from $x_{t}$. The reward model obtained from the previous phase is applied to this denoised image, generating the expected preference score $r_{\theta}(c,x^{\prime}_{0})$. This preference score enables the fine-tuning of the diffusion model to align more closely with human preferences:
$\mathcal{L}(\theta)_{refl}=\mathbb{E}_{c\sim p(c)}\mathbb{E}_{x^{\prime}_{0}\sim p(x^{\prime}_{0}|c)}[-r(x^{\prime}_{0},c)]$ (3)
@@ 4 UniFL: Unified Feedback Learning
Our proposed method, UniFL, aims to improve the stable diffusion in various aspects, including visual generation quality, human aesthetic quality, and inference efficiency. our method takes a unified feedback learning perspective, offering a comprehensive and streamlined solution.
An overview of UniFL is illustrated in Fig.2. In the following subsections, we delve into the details of three key components: perceptual feedback learning to enhance visual generation quality (Sec. 4.1); decoupled feedback learning to improve aesthetic appeal (Sec. 4.2); and adversarial feedback learning to facilitate inference acceleration (Sec. 4.3).
Figure 2: The overview of UniFL, which leverages an unified feedback learning framework to comprehensively enhance the model performance and inference speed. The training process of UniFL is divided into two stages, the first stage aims to improve visual quality and aesthetics, and the second stage speeds up model inference. The gray area within the denoise steps are the timesteps where feedback learning optimizes.
@@ 4.1 Perceptual Feedback Learning
Current diffusion models exhibit limitations in achieving high-quality visual generation, particularly in areas such as image style shift and object structure distortion. These limitations stem from the reliance on reconstruction loss solely in the latent space, which lacks supervision based on visual perception in the image space. To address this issue, as illustrated in Fig.3, we propose Perceptual Feedback Learning (PeFL) to fine-tune diffusion model using visual feedback provided by existing perceptual models. Our key insight is that various visual perception models already encapsulate rich visual priors from diverse aspects. The complete PeFL process is summarized in Algorithm 1. In contrast to ReFL, which starts from a randomly initialized latent representation and only considers the text prompt as a condition, PeFL incorporates image content as an additional visual condition for perceptual guidance. Specifically, given a text-image pair, $(c,x)$, we first select a forward step $T_{a}$ and inject noise into the ground truth image to obtain a conditional latent sequence $x_{0}\rightarrow x_{1}...\rightarrow x_{T_{a}}$. Subsequently, we randomly select a denoising time step $t$ and denoising from $x_{T_{a}}$, yielding $x_{T_{a}}\rightarrow x_{T_{a}-1}...\rightarrow x_{t}$. Next, we directly predict $x_{t}\rightarrow x^{\prime}_{0}$. Various perceptual models are employed to provide feedback on $x^{\prime}_{0}$ for specific visual aspects:
i) Style: To capture image style, we employ the VGG model to encode image features and extract visual style using the widely adopted gram matrix in style transfer. The feedback on style is calculated as follows:
$\mathcal{L}(\theta)_{pefl\_style}=\mathbb{E}_{x_{0}\sim\mathcal{D},x^{\prime}_{0}\sim G(x_{t_{a}})}\|Gram(V(x^{{}^{\prime}}_{0}))-Gram(V(x_{0}))\|_{2},$ (4)
where $V$ is the VGG network, and $Gram$ is the calculation of the gram matrix.
ii) Structure: For extracting visual structure information, we utilize visual instance segmentation models, as instance masks provide fundamental object structure descriptions. The objective is formulated as:
@@
$\mathcal{L}(\theta)_{pefl\_structure}=\mathbb{E}_{x_{0}\sim\mathcal{D},x^{\prime}_{0}\sim G(x_{t_{a}})}\mathcal{L}_{instance}(m_{I}(x^{{}^{\prime}}_{0}),GT(x_{0}))$ (5)
where $m_{I}$ is the instance segmentation model, $GT(x_{0})$ is the ground truth instance segmentation annotation and $\mathcal{L}_{instance}$ is the instance segmentation loss.
The flexibility of PeFL allows us to leverage various existing visual perceptual models, for example, semantic segmentation models, to provide specific visual feedback. More experiments and results can be found in the Appendix.
Algorithm 1 Perceptual Feedback Learning (PeFL) for LDMs

### v3

# UniFL:  Improve Latent Diffusion Model via Unified Feedback Learning

@@ Abstract
[…]
Figure 1: Generated samples with 20 steps inference from stable-diffusion-xl-base-1.0 optimized by Unified Feedback Learning (UniFL). The last three images of the third row are generated with 4 steps.
@@ 1 Introduction
The emergence of diffusion models has led to remarkable advances in the field of text-to-image (T2I) generation, marked by notable milestones like DALLE-3 [1], Imagen [2], Midjourney [3], etc, elevating the generation quality of images to an unprecedented level. Particularly, the introduction of open-source image generation models, exemplified by latent diffusion model (LDM) [4], has inaugurated a transformative era of text-to-image generation, triggering numerous downstream applications such as T2I personalization [5, 6, 7, 8], controllable generation [9, 10, 11] and text-to-video (T2V) generation [12, 13, 14]. Nevertheless, despite these advancements achieved thus far, current latent diffusion-based image generation models still exhibit certain limitations. i) Inferior visual quality: The generated images still suffer from poor visual quality and lack authenticity. Examples include characters with incomplete limbs or distorted body parts, as well as limited fidelity in terms of style representation. ii) Inadequate aesthetic appeal: The generated image tends to lack aesthetic appeal and often fails to align with human preferences, especially in the abstract aesthetic concepts aspects such as color, lighting, atmosphere, etc. iii) Slow inference speed: The iterative denoising process employed by diffusion models led to inefficiencies during inference that significantly impede generation speed, thereby limiting the practicality of these models in various application scenarios. Recently, numerous works have endeavored to address the aforementioned challenges. For instance, RAPHAEL [15] resorts to the techniques of Mixture of Experts) [16, 17, 18] boost the generation performance via stacking the space MoE and time MoE block. Works [19, 20, 21, 22, 23] represented by ImageReward [23] propose incorporating human preference feedback to guide diffusion models toward aligning with human preferences. SDXL Turbo [24], PGD [25], and LCM [26, 27], on the other hand, targets on achieve inference acceleration through techniques like distillation and consistency models [28]. However, these methods primarily concentrate on tackling individual problems through specialized designs, which poses a significant challenge to the elegant integration of these techniques. For example, MoE significantly complicates the pipeline, making the acceleration method infeasible to apply, and the consistency models [28] alter the denoising process of the diffusion model, making it arduous to directly apply the ReFL preference tuning framework proposed by ImageReward [23]. Therefore, a natural question arises: Can we devise a more effective approach that comprehensively enhances diffusion models in terms of image quality, aesthetic appearance, and generation speed?
To tackle this issue, we present UniFL, a solution that offers a comprehensive improvement to latent diffusion models through unified feedback learning formulation. UniFL aims to boost the visual generation quality, enhance aesthetic attractiveness, and accelerate the inference process. To achieve these objectives, UniFL features three novel designs upon the unified formulation of feedback learning. Firstly, we introduce a pioneering perceptual feedback learning (PeFL) framework that effectively harnesses the extensive knowledge embedded within diverse existing perceptual models to provide more precise and targeted feedback on the potential visual defects of the generated results. Secondly, we employ decoupled aesthetic feedback learning to boost the visual appeal, which breaks down the coarse aesthetic concept into distinct aspects such as color, atmosphere, and texture, simplifying the challenge of abstract aesthetic optimization. Furthermore, an active prompt selection strategy is also introduced to choose the more informative and diverse prompt to facilitate more efficient aesthetics preference learning. Lastly, UniFL develops adversarial feedback learning to achieve inference acceleration by incorporating the adversarial objective in feedback tuning. We instantiate UniFL with a two-stage training pipeline and validate its effectiveness with SD1.5 and SDXL, yielding impressive improvements in generation quality and acceleration. Our contributions are summarized as follows:
@@
  • Novelty and Pioneering: In our work, we shed light on the untapped potential of leveraging existing perceptual models in feedback learning for diffusion models. We highlight the significance of decoupled reward models and elucidate the underlying acceleration mechanism through adversarial training.
  • High Effectiveness: Through extensive experiments, we demonstrate the substantial improvements achieved by UniFL across various types of diffusion models, including SD1.5 and SDXL, in terms of generation quality and inference acceleration.
2 Related Works
Text-to-Image Diffusion Models. Text-to-image generation has gained unprecedented attention over other traditional tasks [29, 30, 31, 32, 33]. Recently, diffusion models have gained substantial attention and emerged as the de facto mainstream method for text-to-image generation, surpassing traditional image generative models like GAN [34] and VAE [35]. Numerous related works have been proposed, including GLIDE [36], DALL-E2 [1], Imagen [2], CogView [37] etc.. Among these, Latent Diffusion Models (LDM) [4] extend the diffusion process to the latent space and significantly improve the training and inference efficiency of the diffusion models, opening the door to diverse applications such as controllable generation [9, 10], image editing [11, 38, 39], and image personalization [5, 7, 6] and so on. Even though, current text-to-image diffusion models still have limitations in inferior visual generation quality, deviations from human aesthetic preferences, and inefficient inference. The target of this work is to offer a comprehensive solution to address these issues.
Improvements on Text-to-Image Diffusion Models. Given the aforementioned limitations, researchers have proposed various methods to tackle these issues. Notably,  [40, 15, 41] focuses on improving generation quality through more advanced training strategies. Inspired by the success of reinforcement learning with human feedback (RLHF) [42, 43] in the field of LLM,  [20, 21, 44, 23, 45] explore the incorporation of human feedback to improve image aesthetic quality. On the other hand,  [25, 24, 28, 27, 26] concentrate on acceleration techniques, such as distillation and consistency models [28] to achieve inference acceleration. While these methods have demonstrated their effectiveness in addressing specific challenges, their independent nature makes it challenging to combine them for comprehensive improvements. In contrast, our study unifies the objective of enhancing visual quality, aligning with human aesthetic preferences, and acceleration through the feedback learning framework.
@@ 3 Preliminaries
Latent Diffusion Model. Text-to-image latent diffusion models leverage diffusion modeling to generate high-quality images based on textual prompts, which generate images from Gaussian noise through a gradual denoising process. During pre-training, a sampled image $x$ is first processed by a pre-trained VAE encoder to derive its latent representation $z$. Subsequently, random noise is injected into the latent representation through a forward diffusion process, following a predefined schedule $\{\beta_{t}\}^{T}$. This process can be formulated as $z_{t}=\sqrt{\overline{{\alpha}}_{t}}z+\sqrt{1-\overline{{\alpha}}_{t}}\epsilon$, where $\epsilon\in\mathcal{N}(0,1)$ is the random noise with identical dimension to $z$, $\overline{{\alpha}}_{t}=\prod_{s=1}^{t}\alpha_{s}$ and $\alpha_{t}=1-\beta_{t}$. To achieve the denoising process, a UNet $\epsilon_{\theta}$ is trained to predict the added noise in the forward diffusion process, conditioned on the noised latent and the text prompt $c$. Formally, the optimization objective of the UNet is:
@@
Reward Feedback Learning. Reward feedback learning (ReFL) [23] is a preference fine-tuning framework that aims to improve the diffusion model via human preference feedback. It consists of two phases: (1) Reward Model Training and (2) Preference Fine-tuning. In the Reward Model Training phase, human preference data is collected to train a human preference reward model, which serves as a proxy to provide human preferences. More specifically, considering two candidate generations, denoted as $x_{w}$ (preferred generation) and $x_{l}$ (unpreferred one), the loss function for training the human preference reward model $r_{\theta}$ can be formulated as follows:
$\mathcal{L}_{\mathrm{rm}}(\theta)=-\mathbb{E}_{(c,x_{w},x_{l})\sim\mathcal{D}}[log(\sigma(r_{\theta}(c,x_{w})-r_{\theta}(c,x_{l})))]$ (2)
where $\mathcal{D}$ denotes the collected feedback data, $\sigma(\cdot)$ represents the sigmoid function, and $c$ corresponds to the text prompt. The reward model $r{\theta}$ is optimized to produce a reward score that aligns with human preferences. In the Preference Fine-tuning phase, ReFL begins with an input prompt $c$, initializing a random latent variable $x_{T}$. The latent variable is then progressively denoised until reaching a randomly selected timestep $t$. Then, the denoised image $x^{\prime}_{0}$ is directly predicted from $x_{t}$. The reward model obtained from the previous phase is applied to this denoised image, generating the expected preference score $r_{\theta}(c,x^{\prime}_{0})$. ReFL maximizes such preference scores to make the diffusion model generate images that align more closely with human preferences:
$\mathcal{L}_{\mathrm{refl}}(\theta)=\mathbb{E}_{c\sim p(c)}\mathbb{E}_{x^{\prime}_{0}\sim p(x^{\prime}_{0}|c)}[-r(x^{\prime}_{0},c)]$ (3)
@@ 4 UniFL: Unified Feedback Learning
Our proposed method, UniFL, aims to improve the latent diffusion models in various aspects, including visual generation quality, human aesthetic quality, and inference efficiency. our method takes a unified feedback learning perspective, offering a comprehensive and streamlined solution. An overview of UniFL is illustrated in Fig.2. In the following subsections, we delve into the details of three key components: perceptual feedback learning to enhance visual generation quality (section 4.1); decoupled feedback learning to improve aesthetic appeal (section 4.2); and adversarial feedback learning to facilitate inference acceleration (section 4.3).
Figure 2: Overview of UniFL. We leverage a unified feedback learning framework to enhance the model performance and inference speed comprehensively. The training process of UniFL is divided into two stages, the first stage aims to improve visual quality and aesthetics, and the second stage speeds up model inference.
@@ 4.1 Perceptual Feedback Learning
Current diffusion models exhibit limitations in achieving high-fidelity visual generation, for example, object structure distortion. These limitations stem from the reliance on reconstruction loss(MSE loss) solely in the latent space, which lacks structural supervision on the high-level visual quality. To address this issue, we propose perceptual feedback learning (PeFL). Our key insight is that various visual perception models already embed rich visual priors, which can be exploited to provide feedback for visual generation and fine-tune the diffusion model. The complete PeFL process is summarized in Algorithm 1. In contrast to ReFL, which starts from a randomly initialized latent representation and only considers the text prompt as a condition, PeFL incorporates image content as an additional visual condition for perceptual guidance. Specifically, given a text-image pair, $(c,x)$, we first select a forward step $T_{a}$ and inject noise into the ground truth image to obtain a conditional latent $x_{0}\rightarrow x_{T_{a}}$. Subsequently, we randomly select a denoising time step $t$ and denoising from $x_{T_{a}}$, yielding $x_{T_{a}}\rightarrow x_{T_{a}-1}...\rightarrow x_{t}$. Next, we directly predict $x_{t}\rightarrow x^{\prime}_{0}$. By incorporating the visual condition input, the denoised image is expected to restore the same high-level visual characteristics, such as object structure, and style, which existing perception models can capture. For instance, in the case of object structure, the instance segmentation model can serve as a valuable resource as it provides essential descriptions of object structure through instance masks. Consequently, the feedback on the generation of such visual characteristics on $x^{\prime}_{0}$ can be obtained by comparing it with the ground truth segmentation mask via:
@@
$\mathcal{L}^{\mathrm{struct}}_{\mathrm{pefl}}(\theta)=\mathbb{E}_{x_{0}\sim\mathcal{D},x^{\prime}_{0}\sim G(x_{t_{a}})}\mathcal{L}_{\mathrm{instance}}(m_{I}(x^{{}^{\prime}}_{0}),\mathrm{GT}(x_{0}))$ (4)
where $m_{I}$ is the instance segmentation model, $\mathrm{GT}(x_{0})$ is the ground truth instance segmentation mask and $\mathcal{L}_{\text{instance}}$ is the instance segmentation loss. Note that our PeFL differs from ReFL as indicated by the red font in Algorithm 1. With the visual condition input and perception model, the diffusion model is allowed to get a detailed and focused feedback signal on a specific aspect, instead of the general quality feedback offered by ReFL. Moreover, the flexibility of PeFL allows us to leverage various existing visual perceptual models, more examples can be found in the Appendix A.
4.2 Decoupled Feedback Learning
Decoupled Aesthetic Fine-tuning. Existing text-to-image diffusion models exhibit shortcomings in images that satisfy human aesthetic preferences. While PeFL prioritizes objective visual quality, aesthetic quality is inherently subjective and abstract, requiring human aesthetic feedback to steer the generation process. Despite ImageReward’s attempt to incorporate human aesthetic preferences through a reward model, its performance is hindered by oversimplified modeling that fails to capture the multidimensional nature of human aesthetic preferences. Generally, humans consider the aesthetic attractiveness of an image from various aspects, such as color, lighting, etc, and conflating these aspects without distinguishing during preference tuning would encounter optimization conflicts as evidenced in  [46]. To address this issue, we follow [23] to achieve aesthetic preference tuning but suggest decoupling the various aesthetic aspects when constructing preference reward models. Specifically, we decomposed the general aesthetic concept into representative dimensions and collected the corresponding annotated data, respectively. These dimensions include color, layout, lighting, and detail. Subsequently, we train a separate aesthetic preference reward model for each annotated data according to Eq.2. Finally, we leveraged these reward models for aesthetic preference tuning:
$\mathcal{L}_{\mathrm{aes}}(\theta)=\sum_{d}^{K}\mathbb{E}_{c\sim p(c)}\mathbb{E}_{x^{\prime}_{0}\sim p(x^{\prime}_{0}|c)}[\texttt{ReLU}(\alpha_{d}-r_{d}(x^{\prime}_{0},c))]$ (5)
$r_{d}$ is the aesthetic reward model on $d$ dimension, $d\in\{\text{color},\text{layout},\text{detail},\text{lighting}\}$, $\alpha_{d}$ is the dimension-aware hinge coefficient, and $K$ is the number of fine-grained aesthetic dimension.
Algorithm 1 Perceptual Feedback Learning (PeFL)